# Observability-Aware Control for Quadrotor Formation Flight with Range-only Measurement

### v2

@@ 5.2 Satisfaction of local weak observability under OPC control
This result also connects our OPC to previous work optimizing metrics of the observability matrix [15]. $C_{1},C_{2}$ in Proposition 5.2 are shown by Propositions B.4 and B.7 to be:
@@ 6 Experimental Validation
We evaluate our observability-aware controller on CL-aided quadrotors in two stages. In Section 6.1, we describe our Monte Carlo simulations, modelled after our previous work [6], of an EKF-based CL system configured according to Fig. 3. We evaluate the localization performance of this system when guided by our controller compared to two baseline trajectories. In Section 6.2, we describe our quadrotor flight experiments using the same EKF-based CLS. Collectively, the simulations and experiments assess the performance of observability-aware control under realistic sensing and quadrotor dynamics, with a focus on estimation accuracy and precision.
To compute the STLOG in Eq. (39) in software, we use jax [23] to carry out automatic differentiation and evaluate the Lie derivative of our system defined by Eq. (4) up to r=5r=5101010Python implementation of the OPC can be found inhttps://github.com/FSC_Lab/observability_aware_controller. However, since there is limited upstream support for jax on board computers with the ARM architecture, we cannot compute STLOG in real-time on board. Instead, we use the OPC to compute control inputs online in simulations, but run it offline as a sliding-window trajectory planner to pre-plan trajectories for the quadrotors in experiments.
@@ 6.1 Simulation
An observability-aware trajectory performed with the OPC computing control inputs in real-time.
Figure 4: Trajectories executed by the quadrotor team in simulations; Squares and circles mark starting and end positions, respectively
@@
Parameter Value Notes
@@
Figure 4 shows the trajectories trialed in simulation. Our optimized trajectories in Case 3 exhibit visually distinct behaviors — including periodic orbiting patterns — to deliberately excite weakly observable directions. These motions emerge from the OPC without specific hints or deliberate configuration, reproducing the helical [24] and orbiting [17] behavior of observability-optimal trajectories in previous works.
(a) X-axis position error; Horizontal lines show RMS error
(b) Estimated X-axis uncertainty ($3\sigma$ envelope)
(c) Y-axis position error; Horizontal lines show RMS error
(d) Estimated Y-axis uncertainty ($3\sigma$ envelope)
(e) Z-axis position error; Horizontal lines show RMS error
(f) Estimated Z-axis uncertainty ($3\sigma$ envelope)
Figure 5: Results of Monte Carlo simulations of EKF-based CLS along trajectories generated in test cases 1, 2, and 3 (ours). Position estimation error, RMSE error, and uncertainty ($3\sigma$ envelope) are shown along each axis
Figures 5a to 5f visualize the estimation performance of the EKF-based CLS in simulations. The straight trajectory yields in Case 1 a steadily diverging error and high covariance due to a lack of perception- or observability-aware excitation, in line with long-standing conclusions about the pathologies of uniform straight trajectories [3]. The zigzag trajectory in Case 2 is dynamically rich but estimation-unware. It enabled the CLS to occasionally reduce position estimation error at the cost of strong spikes in error that drove up the RMS error. This confirms that simply exciting position space, adding curvature, or desynchronizing follower motion relative to the leader does not guarantee informative sensor observations. Geometrically rich but non-optimized minimum-snap trajectories may also worsen estimation due to poor alignment with observable modes and increased tracking difficulty. In contrast, the observability-optimal (OO) trajectory explicitly excites low-observability directions and maintains low positioning error and tight estimator confidence bounds.
@@
During the simulation, executing the OPC consumes $0.5566\pm 0.1108s$ (mean $\pm$ std) per loop, requiring an average of 40 NLP solver iterations on average. This cost scales linearly with the number of stages $N$ in the horizon. Although this remains too slow for high-rate feedback control, it approaches acceptable upper margins for sparser online replanning.
@@ 6.2 Quadrotor Flight Experiments
Figure 6: A practical cooperative localization system based on UWB interrobot ranging mounted on a pair of F450 quadrotors, portrayed in flight during our flight test experiments
@@
Type Name and Specifications
Airframe F450 (1.7 kg, 5“ props, 4500 mAh LiPo)
Autopilot Cube Orange (PX4 1.14.5)
Onboard Computer Jetson TX2 NX (JetPack 4.6)
GNSS unit (leader) UBlox M8N
UWB unit (follower) ESP32 UWB DWM3000
@@
We use a range-only CLS interfaced with a UWB range system [^11] The C/C++ interface to the UWB sensor can be found at: https://github.com/FSC_Lab/dwm3000_ros to estimate the states of both quadrotors, so that the estimation performance can be compared between the two test cases.
Figures 7a and 7b show straight (Case 1) and observability-aware (Case 2) trajectories during flight tests. The observability-aware trajectory induces an orbiting motion in the follower, while the leader continues along a uniform straight path. This motion pattern yields significantly more excitation in the inter-vehicle range.
(a) Quadrotor trajectories in flight experiment Case 1.
(b) Quadrotor trajectories in flight experiment Case 2.
Figure 7: Trajectories executed by the quadrotor team in flight experiments Case 1 and 2 (ours)
@@
Table 6: Positioning Uncertainty of EKF-based CLS along different trajectories in flight tests. Position estimation error, RMSE error, and uncertainty ($3\sigma$ envelope) are shown along each axis
@@
In the absence of ground-truth data, the positioning error in Fig. 8 is relative to the postprocessed GNSS measurements, which have a positioning variance of approximately 0.3 m. However, this error is still broadly illustrative of the CLS performance. Our optimized trajectory in Case 2 maintained stable and bounded positioning error on both the X and the Y axes. In contrast, the straight trajectory in Case 1 exhibited asymmetric performance: better accuracy on the X-axis but significantly worse on the Y-axis, reflecting poor excitation in that dimension. This shows that the optimized trajectory guarantees an acceptable positioning performance evenly across all axes, which is consistent with our goal of maximizing the minimum improvement in precision along the least observable directions.
@@
(a) X-axis position error
(b) Estimated X-axis uncertainty ($3\sigma$ envelope)
(c) Y-axis position error
(d) Estimated Y-axis uncertainty ($3\sigma$ envelope)
(e) Z-axis position error
(f) Estimated Z-axis uncertainty ($3\sigma$ envelope)
Figure 8: Position estimation performance of EKF-based CLS during flight experiments Case 1 and 2 (ours); Squares and circles mark starting and end positions, respectively
@@ 7 Conclusion
We proposed an observability-aware controller to improve positioning precision for quadrotor drones that perform formation flight using range-only CL. This controller addresses the issue that a part of the system state is not directly observable through range measurements, and higher estimation precision can be achieved by a control strategy that maximizes observability. This controller builds on recent work on the LOG, which captures information about the system’s observability better than first-order quantities such as the FIM, CRLB, or the EKF covariance.
We derived STLOG, a novel point-wise approximation of the LOG. Using the STLOG, we formulate an observability-optimal control principle that maximizes the sum of the minimum improvements in estimation precision along a trajectory. We designed the OPC, a receding-horizon controller, based on this principle. The OPC refines the estimation in the least observable directions of the quadrotor state and improves the reliability of range-only CL.
We validated our approach using the accuracy and precision of the estimation as performance metrics when we applied the OPC to a range-only quadrotor CLS. Our Monte Carlo simulations showed consistent improvements in estimator performance along OPC-generated trajectories compared to both straight and artificially perturbed trajectories. These improvements were further supported by flight experiments, in which our OPC was deployed on a quadrotor team performing CL.
Appendix A Formulas regarding likelihood ratios of infinite-dimensional spaces
In Section 4, we required the computation of likelihood ratios of the form $p(\mathbf{y}(t)|\mathbf{x}_{0})/p(\mathbf{y}(t))$ (e.g. in Eq. (16)). These likelihood ratios can be formalized as Radon–Nikodym densities of the underlying probability measures on infinite-dimensional spaces. In this appendix we outline the construction of these measures using Gaussian measure theory in general. We assume familiarity with the measure-theoretic formulation of probability in this section, for which we refer to [25]. We also use established results in Gaussian measure theory extensively, for which we refer to [26, 27].
First, we state the main results of Gaussian measure theory on separable Hilbert spaces. Let $H$ be a separable Hilbert space with inner product $\langle\cdot,\cdot\rangle_{H}$. A Borel probability measure $\mu$ on $H$ is called a Gaussian measure if, for every continuous linear functional $l\in H^{*}$, the pushforward measure $l_{*}\mu$ on $\mathbb{R}$ is a one-dimensional Gaussian measure [^12] Here we define a one-dimensional Gaussian measure on $\mathbb{R}$ as a Borel measure $\mu$ whose Fourier transform is given as $\int_{\mathbb{R}}e^{-ikx}d\mu(x)=\exp(imk-qk^{2})$ for some real parameters $m\in\mathbb{R},q\in[0,+\infty)$, so it includes the Dirac measure in particular.. Moreover, we call the measure $\mu$ centered if $l_{*}\mu$ has zero mean for all $l\in H^{*}$.
The mean $m:H^{*}\rightarrow\mathbb{R}$ and covariance $\mathcal{K}:H^{*}\times H^{*}\rightarrow\mathbb{R}$ of the Gaussian measure $\mu$ is then defined as the continuous linear (resp. bilinear) functionals
$m(l)=\int_{H}l(y)\ d\mu(y),\quad\mathcal{K}(l,l^{\prime})=\int_{H}l(y-m)l^{\prime}(y-m)\ d\mu(y),$ (44)
for all $l,l^{\prime}\in H^{*}$. Here we identified $m\in H^{**}\cong H$ in Eq. (44). Moreover, the covariance $\mathcal{K}$ in Eq. (44) induces a bounded, self-adjoint and trace-class operator on $H$ through the canonical isomorphisms $H\cong H^{*}\cong H^{**}$, which we also denote as $\mathcal{K}$ by abuse of notation.
Conversely, let $\mathcal{K}:H\rightarrow H$ be a bounded, self-adjoint and trace-class operator. Then $\mathcal{K}$ defines a centered Gaussian measure $\mu_{0,\mathcal{K}}$ using the formulas in Eq. (44) with $m=0$. Consider the canonical map $\iota:H^{*}\rightarrow L^{2}(H,\mu_{0,\mathcal{K}})$, where we consider a linear functional of $H$ as a $\mu_{0,\mathcal{K}}$-square-integrable function. The closure of $\iota(H^{*})$ in $L^{2}(H,\mu_{0,\mathcal{K}})$ under the Hilbert space norm is known as the reproducing kernel Hilbert space (RKHS) $\mathcal{R}_{\mathcal{K}}$, and is itself a Hilbert space with the inner product
$\langle\iota(l),\iota(l^{\prime})\rangle=\int_{H}l(y)l^{\prime}(y)\ d\mu_{0,\mathcal{K}}(y)=\mathcal{K}(l,l^{\prime}).$ (45)
[…]
@@ Appendix A Formulas regarding likelihood ratios of infinite-dimensional spaces
and we can show in fact that $h\mapsto\hat{h}$ is an isomorphism of Hilber spaces $\mathcal{H}_{\mathcal{K}}\cong\mathcal{R}_{\mathcal{K}}$ when the Cameron–Martin space is equipped with the following inner product:
@@
The Cameron–Martin theorem characterizes when two Gaussian measures with the same covariance are absolutely continuous w.r.t. each other:
@@
The Gaussian measure $\mu_{h,\mathcal{K}}$ is absolutely continuous with respect to $\mu_{0,\mathcal{K}}$, if and only if $h\in\mathcal{H}_{\mathcal{K}}$. In this case, the Radon–Nikodym density is
$\frac{d\mu_{h,\mathcal{K}}}{d\mu_{0,\mathcal{K}}}(y)=\exp\left(\hat{h}(y)-\frac{1}{2}\langle h,h\rangle_{\mathcal{H}_{K}}\right),$ (48)
where $\hat{h}$ and $\langle h,h\rangle_{\mathcal{H}_{\mathcal{K}}}$ are given in Eqs. (46) and (47), respectively.
Now consider the spectral decomposition of $\mathcal{K}$: let ${\{e_{i}\}}_{i=1}^{\infty}$ be an orthonormal eigenbasis for $\mathcal{K}$ with corresponding eigenvalues $\{\lambda_{i}\}_{i=1}^{\infty}$. If $\lambda_{i}>0$ for all $i$, we have the following expedient characterization:
@@
On the contrary, if $\mathcal{K}$ only has a finite number of non-zero eigenvalues, then $\mathcal{H}_{\mathcal{K}}=\text{range}(\mathcal{K})$ algebraically, and in particular we can only compute likelihood ratios when we constrain the mean in a finite-dimensional subspace.
In the following, we focus on the case where $\mathcal{K}$ is positive definite i.e. has no zero eigenvalues.
The Radon–Nikodym density in Eq. (48) is merely a $\mu_{0,\mathcal{K}}$-measurable function on $H$ for a fixed $h\in\mathcal{H}_{\mathcal{K}}$. An easy way to obtain more uniformity on both $h$ and $y\in H$ is to only consider mean displacements in a smaller subspace $\mathcal{H}_{\mathcal{K}^{2}}\subset\mathcal{H}_{\mathcal{K}}$. Since $\mathcal{H}_{\mathcal{K}^{2}}\hookrightarrow\mathcal{H}_{\mathcal{K}}\hookrightarrow H$ and $\mathcal{H}_{\mathcal{K}^{2}}\xrightarrow[]{\mathcal{\mathcal{K}}^{-1/2}}\mathcal{H}_{\mathcal{K}}\xrightarrow[]{\mathcal{K}^{-1/2}}H$ are continuous, we have:
@@
Let $\mathcal{K}$ be a positive definite trace-class operator on a separable Hilbert space $H$, and let $\mathcal{H}_{\mathcal{K}^{2}}$ be the Cameron–Martin space of $\mathcal{K}^{2}$. For $h\in\mathcal{H}_{\mathcal{K}^{2}}$, the Radon–Nikodym density of $\mu_{h,\mathcal{K}}$ with respect to $\mu_{0,\mathcal{K}}$ exists, and moreover is a continuous function on $\mathcal{H}_{K^{2}}\times H$ given by
@@
where $F$ is the positive function given by Eq. (49) and $\otimes$ denotes the product measure. Since $(x,y)\mapsto F(\mathcal{A}(x),y)$ is continuous, it is Borel measurable and $\nu\otimes\mu_{0,\mathcal{K}}$-integrable. We check by Tonelli’s theorem that
@@
Using the Cameron–Martin formula in Eq. (48) in the penultimate step, we can verify that the conditional measure $\gamma|_{x}$ of $\gamma$ in Eq. (50) at a given $x\in X$ and the marginal measure $\bar{\gamma}$, both measures on $H$, are given respectively by
$\gamma|_{x}(y)=F(\mathcal{A}(x),y)\mu_{0,\mathcal{K}}(y)=\mu_{\mathcal{A}(x),\mathcal{K}}(y),\quad\bar{\gamma}(y)=\left(\int_{X}F(\mathcal{A}(x),y)\ d\nu(x)\right)\mu_{0,K}(y),$ (51)
from which it follows that both $\gamma|_{x}$ and $\bar{\gamma}$ have positive Radon–Nikodym densities with respect to $\mu_{0,\mathcal{K}}$, and therefore are absolutely continuous with respect to each other. We summarize our findings in Proposition A.2:
@@
A useful special case is when $H=L^{2}([0,T],\mathbb{R}^{p})$, the Hilbert space of square-intergrable $\mathbb{R}^{p}$-valued functions on the closed interval $[0,T]$. Our convention for the Hilbert space inner product is
@@
The continuity of the map $\mathcal{A}:X\rightarrow\mathcal{H}_{\mathcal{K}^{2}}$ required by Proposition A.2 is often more expediently checked by introducing some intermediate Sobolev space $H^{s}([0,T],\mathbb{R}^{p})$.
@@
    $\bar{\gamma}$, the marginal distribution on $L^{2}([0,T],\mathbb{R}^{p})$ obtained from marginializing $x$ in the joint distribution in Eq. (50),
@@
Simply note that the condition in Eq. (54) is equivalent to the continuity of the embedding $\iota:H^{s}([0,T],\mathbb{R}^{p})\hookrightarrow\mathcal{H}_{\mathcal{K}^{2}}$. Then $\mathcal{A}:X\rightarrow H^{s}([0,T],\mathbb{R}^{p})\hookrightarrow\mathcal{H}_{\mathcal{K}^{2}}$ is continuous, and the assumptions of Proposition A.2 are met. ∎
@@
Condition 1 of Corollary A.2.1 is often the ‘best-case’ result we can hope for if $\mathcal{A}$ is a map that involves the solution of ordinary differential equations. In practice, we can assume that $\mathcal{A}:X\rightarrow C^{\infty}([0,T],\mathbb{R}^{p})$ is continuous with respect to the Fréchet topology in the image, so in particular the map is continuous into Sobolev spaces of all orders.
Example A.1.
We give a recipe for constructing covariance operators $\mathcal{K}$ satisfying condition 2 of Corollary A.2.1. We discuss only the case for scalar-valued functions, i.e. $p=1$. The vector-valued case follows easily.
Let ${\left\{e_{n}(t)\right\}}_{n=0}^{\infty}$ be the orthonormal Legendre polynomials on $[0,T]$ under the inner product in Eq. (53). For a given function $h\in L^{2}([0,T])$, let $h_{n}=\langle h,e_{n}\rangle_{L^{2}([0,T])}$ be its $n$-th Fourier–Legendre coefficient. From [28], inequality (5.4.12), it follows that

### v3

@@ 5.2 Satisfaction of local weak observability under OPC control
This result also connects our OPC to previous work optimizing metrics of the observability matrix [sharma2014observability]. $C_{1},C_{2}$ in Proposition 5.2 are shown by Propositions B.4 and B.7 to be:
@@ 6 Simulation and Experimental Validation
We evaluate our OPC on CLS-aided quadrotors in two stages. In Section 6.1, we describe our Monte Carlo simulations, modelled after our previous work [go2024trajectory], of an EKF-based CLS configured according to Fig. 4. We evaluate the localization performance of this system when guided by our controller compared to two baseline trajectories. In Section 6.2, we describe our quadrotor flight experiments using the same EKF-based CLS.
The computational cost of the OPC is dominated by two factors: evaluating the STLOG and solving the receding-horizon optimal control problem. To evaluate the STLOG, we compute Lie derivatives by automatic differentiation (via jax [jax2018github] [^10] Python implementation of the OPC can be found inhttps://github.com/FSC-Lab/observability_aware_controller), which is competitive with quadratic Lie-derivative approximations used in Ref. [yu2015observability] while avoiding truncation errors and simplifying implementation. Representative evaluation and solve times are summarized in Table 3.
Table 3: STLOG evaluation time
Component Setting Typical time Worst-case time
STLOG evaluation $r=5$ 0.05–0.15 s << 0.3 s
Objective evaluation $r=5$, $N=20$ 0.1–0.3 s << 1.0 s
Total OPC solve $r=5$, $N=20$, 40 iters max 2.0–4.0 s << 6.0 s
The optimization cost scales primarily with the horizon length $N$ and the number of solver iterations. With the settings reported in Table 3, solve times are compatible with online, receding-horizon execution at guidance-level update rates, but are not yet suitable for real-time thrust and body-rate generation on embedded hardware. Therefore, online execution of the OPC is evaluated in simulation, while in experiments, the OPC is used as an offline trajectory generator.
Figure 5: Simulated trajectories with squares/circles marking starting/end positions, respectively.
@@ 6.1 Simulation
An observability-aware trajectory performed with the OPC computing control inputs online.
@@
Noise Type Parameter Value Notes
@@
Figure 5 shows the trajectories trialed in Monte Carlo simulations. Our optimized trajectories in Case 3 exhibit distinct behaviors to deliberately excite weakly observable directions. This periodic dynamic formation emerges from the OPC without specific hints or deliberate configuration, reproducing the helical [li2022three] and orbiting [boyinine2022observability] behavior of observability-optimal trajectories in previous works. Furthermore, our optimized trajectories implied moderate accelerations, as summarized in Table 5, compatible with quadrotor dynamics and control.
Table 5: Acceleration implied by different trajectories in simulations
Trajectory Peak X Acceleration, $m/s^{2}$ Peak Y Acceleration, $m/s^{2}$ Peak Z Acceleration, $m/s^{2}$
Case 3 (ours) 1.737 2.063 1.644
(a) X-axis position error; Horizontal lines show RMS error.
(b) Estimated X-axis uncertainty ($3\sigma$ envelope).
(c) Y-axis position error; Horizontal lines show RMS error.
(d) Estimated Y-axis uncertainty ($3\sigma$ envelope).
(e) Z-axis position error; Horizontal lines show RMS error.
(f) Estimated Z-axis uncertainty ($3\sigma$ envelope).
Figure 6: CLS performance in simulations, quantified by per-axis estimation error, RMSE, and uncertainty.
Figures 6a to 6f show the CLS performance in simulations. The straight trajectory in Case 1 yields a steadily diverging error and high covariance due to a lack of observability-aware excitation, in line with long-standing conclusions about the pathologies of uniform formations [trawny2004optimized]. The zigzag trajectory in Case 2 is dynamically rich but estimation-unaware. It enabled the CLS to occasionally reduce position estimation error at the cost of strong spikes in error that drove up the RMS error. This confirms that simply exciting position space or desynchronizing follower motion relative to the leader does not guarantee informative sensor observations. In contrast, the observability-optimal dynamic formation explicitly excites low-observability directions and maintains low positioning error and tight estimator confidence bounds.
@@ 6.2 Quadrotor Flight Experiments
Figure 7: A UWB-ranging based CLS on a pair of F450 quadrotors during flight tests
@@
Hardware Name and Specifications Software Name/Specifications
Airframe F450 (1.7kg with LiPo, 10“ props) SDK Framework mavros
Autopilot Cube Orange (PX4 1.14.5) Setpoint Type mavros_msgs::AttitudeTarget1
Onboard Computer Jetson TX2 NX (JetPack 4.6) Setpoint Rate Upsampled to 50 Hz
GNSS unit (leader) UBlox M8N Leader Controller Geometric tracking controller
UWB unit (follower) ESP32 UWB DWM30002 Leader Estimator GPS-driven EKF
@@
We use a range-only CLS interfaced with a UWB range system to estimate the states of both quadrotors, so that the estimation performance can be compared between the two test cases.
Figures 8a and 8b show straight (Case 1) and observability-aware (Case 2) trajectories during flight tests. The observability-aware trajectory induces an orbiting motion in the follower, while the leader continues along a uniform straight path. This motion pattern yields significantly more excitation in the inter-vehicle range.
(a) Quadrotor trajectories in flight experiment Case 1..
(b) Quadrotor trajectories in flight experiment Case 2..
Figure 8: Experimental trajectories with squares/circles marking starting/end positions, respectively.
@@
Table 9: CLS performance in flight, quantified by per-axis estimation error, RMSE, and uncertainty
@@
[…]
@@
(a) X-axis position error.
(b) Estimated X-axis uncertainty ($3\sigma$ envelope).
(c) Y-axis position error.
(d) Estimated Y-axis uncertainty ($3\sigma$ envelope).
(e) Z-axis position error.
(f) Estimated Z-axis uncertainty ($3\sigma$ envelope).
Figure 9: CLS positioning performance in flight, with squares/circles marking starting/end positions, respectively.
@@ 7 Conclusion
This work demonstrates that observability-aware control enables robust range-only cooperative localization systems (CLS) for quadrotors. We addressed the fragility of range-only relative localization, where specific state components become poorly observable without sufficiently rich motion.
The proposed STLOG builds upon the classical LOG, which captures information lost by first-order quantities like the FIM, and enables tractable, point-wise evaluation of observability along a trajectory. The resulting OPC serves as both a robust controller and safety mechanism: by maximizing the minimum eigenvalue of the STLOG, it generates motions that excite poorly observable directions, guarantees a minimum precision improvement in state estimation over each control stage, and prevents catastrophic uncertainty growth and estimator divergence.
Our simulations and experiments confirm that observability-aware dynamic formations outperform rigid and perturbed baselines in terms of estimation precision. These results validated the proposed framework for application to quadrotor teams, specifically for leader-follower ferrying missions in GNSS-denied environments. Our leader-follower configuration is limited to two quadrotors, for which a centralized cooperative localization and control architecture is sufficient for simplicity, but our inter-drone dynamics model can be generalized beyond quadrotors to pairs of multirotors in general. The richness of this model incurs a high computational cost in evaluating its observability, which currently restricts the OPC to an offline guidance planner role in experiments. We leave the extension of our approach to distributed estimation and control, and addressing computational bottlenecks, to future work.
Appendix A Derivation of closed form posterior uncertainty
This appendix supports the derivation of the likelihood ratio $p\left(\frac{\mathbf{y}(t)|\mathbf{x}_{0}}{\mathbf{y}(t)}\right)$ proposed in Eq. (19) within Sec. 4 to quantify posterior uncertainty through continuous-time measurements. Likelihood ratios in this form are Radon–Nikodym densities between probability measures on an infinite-dimensional path space. In the Gaussian setting, they admit a closed form via the Cameron–Martin theorem. We refer to [pollard2002user] for background on measure-theoretic probability, and to [bogachev2015gaussian, hairer2023introductionstochasticpdes] on Gaussian measure theory.
We recall the Cameron–Martin theorem, together with the definitions of the Reproducing Kernel Hilbert Space (RKHS), the Cameron–Martin space, and the associated Riesz representations. The Cameron–Martin theorem characterizes when two Gaussian measures with the same covariance are absolutely continuous w.r.t. each other.
Let $H$ be a separable Hilbert space with inner product $\langle\cdot,\cdot\rangle_{H}$. Let $\mu$ be a centered Gaussian measure on $H$, meaning that for every continuous linear functional $l\in H^{*}$, the pushforward measure $l_{*}\mu$ on $\mathbb{R}$ is a one-dimensional Gaussian measure with zero mean. The mean $m:H^{*}\rightarrow\mathbb{R}$ and covariance $\mathcal{K}:H^{*}\times H^{*}\rightarrow\mathbb{R}$ of $\mu$ are defined as the continuous linear functionals $m(l)=\int_{H}l(y)\ d\mu(y)$ and $\mathcal{K}(l,l^{\prime})=\int_{H}l(y-m)l^{\prime}(y-m)\ d\mu(y)$ for all $l,l^{\prime}\in H^{*}$.
In the following, we focus on the covariance operator $\mathcal{K}:H\rightarrow H$. By the Riesz identification $H\simeq H^{*}$, we use the same symbol $\mathcal{K}$ for the induced bilinear form on $H^{*}\times H^{*}$. We assume $\mathcal{K}$ is a bounded, self-adjoint, and trace-class operator. Then, there exists a centered Gaussian measure $\mu_{0,\mathcal{K}}$ with covariance $\mathcal{K}$. All subsequent constructions are expressed in terms of $\mathcal{K}$.
Definition A.1 (Reproducing kernel Hilbert space).
Let $\iota:H^{*}\to L^{2}(H,\mu_{0,\mathcal{K}})$ denote the canonical embedding that identifies linear functionals with square-integrable functions. The closure of $\iota(H^{*})$ in $L^{2}(H,\mu_{0,\mathcal{K}})$ under the Hilbert space norm is the reproducing kernel Hilbert space (RKHS) $\mathcal{R}_{\mathcal{K}}$, and is itself a Hilbert space with the inner product $\langle\iota(l),\iota(l^{\prime})\rangle=\int_{H}l(y)l^{\prime}(y)\ d\mu_{0,\mathcal{K}}(y)=\mathcal{K}(l,l^{\prime})$.
Definition A.2 (Cameron–Martin space).
The Cameron–Martin space $\mathcal{H}_{\mathcal{K}}$ is the subset of $H$ defined by $\mathcal{H}_{\mathcal{K}}=\left\{h\in H:\sup\{l(h):l\in H^{*},\mathcal{K}(l,l)\leq 1\}<\infty\right\}$.
Definition A.3 (Riesz representation in the Cameron–Martin space).
For $h\in\mathcal{H}_{\mathcal{K}}$, the Riesz representation $\hat{h}\in\mathcal{R}_{\mathcal{K}}$ is defined by:
@@ Appendix A Derivation of closed form posterior uncertainty
and $h\mapsto\hat{h}$ is an isomorphism of Hilbert spaces $\mathcal{H}_{\mathcal{K}}\cong\mathcal{R}_{\mathcal{K}}$ when the Cameron–Martin space is equipped with the following inner product:
@@
We recall the Cameron–Martin theorem below.
@@
The Gaussian measure $\mu_{h,\mathcal{K}}$ is absolutely continuous with respect to $\mu_{0,\mathcal{K}}$, if and only if $h\in\mathcal{H}_{\mathcal{K}}$. In this case, the Radon–Nikodym density is:
$\frac{d\mu_{h,\mathcal{K}}}{d\mu_{0,\mathcal{K}}}(y)=\exp\left(\hat{h}(y)-\frac{1}{2}\langle h,h\rangle_{\mathcal{H}_{\mathcal{K}}}\right),$ (46)
Let ${\{e_{i}\}}_{i=1}^{\infty}$ be an orthonormal eigenbasis for $\mathcal{K}$ with corresponding eigenvalues $\{\lambda_{i}\}_{i=1}^{\infty}$. If $\lambda_{i}>0$ for all $i$, the following characterization holds:
@@
To obtain a likelihood ratio that depends continuously on both the observation trajectory and the finite-dimensional quantity $\mathbf{x}_{0}$, we focus on the case where $\mathcal{K}$ is positive definite, and only consider mean displacements in a smaller subspace $\mathcal{H}_{\mathcal{K}^{2}}\subset\mathcal{H}_{\mathcal{K}}$. In the context of Sec. 4, $\mathbf{x}_{0}\in\mathcal{X}$ corresponds to the initial condition, while $\mathcal{A}(\mathbf{x})$ denotes the induced observation trajectory. Since $\mathcal{H}_{\mathcal{K}^{2}}\hookrightarrow\mathcal{H}_{\mathcal{K}}\hookrightarrow H$ and $\mathcal{H}_{\mathcal{K}^{2}}\xrightarrow[]{\mathcal{\mathcal{K}}^{-1/2}}\mathcal{H}_{\mathcal{K}}\xrightarrow[]{\mathcal{K}^{-1/2}}H$ are continuous, we have:
@@
Let $\mathcal{K}$ be a positive definite trace-class operator on a separable Hilbert space $H$, and let $\mathcal{H}_{\mathcal{K}^{2}}$ be the Cameron–Martin space of $\mathcal{K}^{2}$. For $h\in\mathcal{H}_{\mathcal{K}^{2}}$, the Radon–Nikodym density of $\mu_{h,\mathcal{K}}$ with respect to $\mu_{0,\mathcal{K}}$ exists, and moreover is a continuous function on $\mathcal{H}_{K^{2}}\times H$ given by:
@@
where $F$ is the positive function given by Eq. (47) and $\otimes$ denotes the product measure. Since $(x,y)\mapsto F(\mathcal{A}(x),y)$ is continuous, it is Borel measurable and $\nu\otimes\mu_{0,\mathcal{K}}$-integrable.
By Tonelli’s theorem, the following normalization ensures that $\gamma$ defines a valid joint distribution compatible with Bayes’ rule in Eq. (19).
@@
Using the Cameron–Martin formula in Eq. (46) in the penultimate step, we verify that the conditional measure $\gamma|_{x}$ of $\gamma$ in Eq. (48) at a given $x\in X$ and the marginal measure $\bar{\gamma}$, both measures on $H$, are given respectively by:
$\gamma|_{x}(y)=F(\mathcal{A}(x),y)\mu_{0,\mathcal{K}}(y)=\mu_{\mathcal{A}(x),\mathcal{K}}(y),\quad\bar{\gamma}(y)=\left(\int_{X}F(\mathcal{A}(x),y)\ d\nu(x)\right)\mu_{0,\mathcal{K}}(y),$ (49)
from which it follows that both $\gamma|_{x}$ and $\bar{\gamma}$ have positive Radon–Nikodym densities with respect to $\mu_{0,\mathcal{K}}$, and therefore are absolutely continuous with respect to each other. The following result provides the likelihood ratio between the conditional and marginal observation measures required by Eq. (19).
@@
A useful special case is when $H=L^{2}([0,T],\mathbb{R}^{p})$, the Hilbert space of square-integrable $\mathbb{R}^{p}$-valued functions on the closed interval $[0,T]$. Our convention for the Hilbert space inner product is
@@
To obtain sufficient conditions in the continuous-time trajectory space used in this work, we specialize Proposition A.2 to an intermediate Sobolev space $H^{s}([0,T],\mathbb{R}^{p})$.
@@
    $\bar{\gamma}$, the marginal distribution on $L^{2}([0,T],\mathbb{R}^{p})$ obtained from marginalizing $x$ in the joint distribution in Eq. (48),
@@
The condition in Eq. (52) is equivalent to the continuity of the embedding $\iota:H^{s}([0,T],\mathbb{R}^{p})\hookrightarrow\mathcal{H}_{\mathcal{K}^{2}}$. Then $\mathcal{A}:X\rightarrow H^{s}([0,T],\mathbb{R}^{p})\hookrightarrow\mathcal{H}_{\mathcal{K}^{2}}$ is continuous, and the assumptions of Proposition A.2 are met. ∎
Finally, we construct a covariance operator $\mathcal{K}$ for scalar-valued functions, satisfying condition 2 of Corollary A.2.1, as a base case that is generalizable to vector-valued functions such as our observation trajectories in Sec. 4.
@@
Let ${\left\{e_{n}(t)\right\}}_{n=0}^{\infty}$ be the orthonormal Legendre polynomials on $[0,T]$ under the inner product in Eq. (51). For a given function $h\in L^{2}([0,T])$, let $h_{n}=\langle h,e_{n}\rangle_{L^{2}([0,T])}$ be its $n$-th Fourier–Legendre coefficient. From [canuto2007spectral], inequality (5.4.12), it follows that: